OpenStudy (anonymous):

Solve the equation: -2sin^2(x)+5cos(x)+1=0, where 0°<=x<360°

OpenStudy (anonymous):

I manipulated it to \[-2(1-\cos^2x)+5cosx+1=0\] is this correct?

OpenStudy (jessereyna):

YAh it looks to be right

OpenStudy (anonymous):

after that i get to \[2\cos^2x+5cosx-1\] and I'm stuck

OpenStudy (landon34):

Simplifying -2sin2(x) + 5cos(x) + 1 = 0 Multiply in2s * x -2in2sx + 5cos(x) + 1 = 0 Multiply cos * x -2in2sx + 5cosx + 1 = 0 Reorder the terms: 1 + 5cosx + -2in2sx = 0 Solving 1 + 5cosx + -2in2sx = 0 Solving for variable 'c'. Move all terms containing c to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + 5cosx + -1 + -2in2sx = 0 + -1 Reorder the terms: 1 + -1 + 5cosx + -2in2sx = 0 + -1 Combine like terms: 1 + -1 = 0 0 + 5cosx + -2in2sx = 0 + -1 5cosx + -2in2sx = 0 + -1 Combine like terms: 0 + -1 = -1 5cosx + -2in2sx = -1 Add '2in2sx' to each side of the equation. 5cosx + -2in2sx + 2in2sx = -1 + 2in2sx Combine like terms: -2in2sx + 2in2sx = 0 5cosx + 0 = -1 + 2in2sx 5cosx = -1 + 2in2sx Divide each side by '5osx'. c = -0.2o-1s-1x-1 + 0.4in2o-1 Simplifying c = -0.2o-1s-1x-1 + 0.4in2o-1 Reorder the terms: c = 0.4in2o-1 + -0.2o-1s-1x-1

OpenStudy (landon34):

i typed this myself. your welcome

OpenStudy (anonymous):

uh they arent variables they're trig functions

OpenStudy (landon34):

I know.

OpenStudy (landon34):

oh, wait.

OpenStudy (landon34):

oops.